OpenStudy (anonymous):

In the equation 234Th90 -> 234Pa91 + X, which particle is represented by X?

OpenStudy (jfraser):

nuclear reactions have to obey conservation of mass just like "regular" chemical reactions. The sum of the masses must remain constant, and the sum of the protons must also remain constant.\[^{234}_{90}Th \rightarrow ^{234}_{91}Pa + ^{A}_{Z}X\] \[234 = 234 + A\] and \[90 = 91 + Z\] the only particle that fits these requirements is the beta particle\[^{0}_{-1} \beta\]

OpenStudy (anonymous):

This is mulitple choice and theres only... a. 0e-1 b. 4He2 c. 1H1 d. 0e+1

OpenStudy (anonymous):

so it wud be A? \[e_{-1}^{0}\]

OpenStudy (xishem):

Yes. That is an electron, which is the product of beta decay.

OpenStudy (anonymous):

Thank You!